OpenStudy (anonymous):

Find the equation in standard form which passes through (-2, -7) and (1, -1). You must show all work

OpenStudy (anonymous):

do you know how to find the slope?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok once you do find the slope just plug it in to the point slope form y-y1=m(x-x1) where: m=slope (x1,y1) -any point at the given line.

OpenStudy (anonymous):

From there you can leave it as it is or solve it for y in order to get y-int form.